# Conversational SimulMT: Efficient Simultaneous Translation with Large Language Models

### v3

@@ 6 Conclusion
Type of LLMs
In this paper, we only experimented with Llama2-7b-chat, which may limit our understanding of the effectiveness of our approach. A more comprehensive study with different-sized LLMs and prompt formats should be conducted.
@@
Our evaluation was only conducted on two commonly used benchmarks which may limit the diversity in domain, style, and language. There may also be a data contamination problem in the experiment. We believe that more datasets should be used to allow for a more comprehensive evaluation.
@@
Appendix A Related Works
Simultaneous Machine Translation (SimulMT)
is the task to provide real-time translation of a source sentence stream where the goal is to minimize the latency while maximizing the translation quality. A common approach is to train a MT model on prefix-to-prefix dataset to directly predict target tokens based on partial source tokens (Ma et al., 2019). Alternatively, Liu et al. (2020) proposed the incremental decoding framework to leverage the pretrained offline NMT and turn it into a SimulMT model without further training. A core component of SimulMT is a read-write policy to decide at every step whether to wait for another source token (READ) or to generate a target token (WRITE). Previous methods have explored fixed policy, which always waits for $k$ tokens before generation  (Ma et al., 2019; Zhang et al., 2022) and adaptive policy, which trains an agent via reinforcement learning (Gu et al., 2017; Arthur et al., 2021). Re-translation (Arivazhagan et al., 2019) from the beginning of the source sentence at the WRITE step will incur high translation latency. Stable hypothesis detection methods such as Local Agreement, hold-$n$ (Liu et al., 2020) and Share prefix SP-n (Nguyen et al., 2021) are employed to commit stable hypothesis and only regenerate a subsequence of source sentence. The goal is to reduce the latency and minimize the potential for errors resulting from incomplete source sentence (Polák et al., 2022; Wang et al., 2021).
LLM-based NMT
Recent research has delved into the potential usage of LLMs in MT Hendy et al. (2023); Zhu et al. (2023); Robinson et al. (2023), especially in handling discourse phenomena (Wang et al., 2023a; Wu et al., 2024) and linguistic nuances such as idioms (Manakhimova et al., 2023). While LLMs do exhibit some level of translation capability, prior research has identified that they still lags behind the conventional NMT models, especially for low resource languages (Robinson et al., 2023). Additionally, the translation performance varies depending on prompting strategies (Zhang et al., 2023). Efforts have been made to enhance the translation performance of LLMs by incorporating guidance from dictionary (Lu et al., 2023), further fine-tuning (Zeng et al., 2023; Xu et al., 2023) and augmenting with translation memories (Mu et al., 2023).
LLM-based SimulMT
SimulLLM (Agostinelli et al., 2023) explore the ability to adapt an LLM finetuned on NMT task to simultaneous translation with wait-k strategy. Wang et al. (2023b) adopt hybrid READ/WRITE policy with wait-k and incremental decoding. TransLLaMA (Koshkin et al., 2024) teach LLMs to produce WAIT tokens to preserve the causal alignment between source and target tokens. At each inference round, LLMs only produce a single word or WAIT token, which is very costly due to multiple rounds of LLM calls. Guo et al. (2024) introduce LLM into the SimulMT task as a translation agent working with a specialized SimulMT policy agent. An additional memory module stores translation history. The policy agent decides on READ/WRITE actions, while the LLM translates target segments. They face the same KV-cache reuse challenge noted by Wang et al. (2023b), making the computational cost of collaborating big and small models even more significant.
Appendix B Simultaneous Decoding with LLM
@@
[…]
@@
Algorithm 1 presents the detail of applying conversational prompt for decoding.

### v4

@@ 7 Conclusion
Our evaluation was conducted on three commonly used benchmarks which may limit the diversity in domains, styles, and languages. There may also be potential data contamination concerns since LLMs might have been exposed to parts of our test sets during pre-training. A more comprehensive evaluation with diverse datasets across more domains and language pairs would strengthen our findings.
Alignment-based Data Curation
Our approach relies on word alignment tools like fast_align to segment parallel sentences, which has inherent limitations. These tools may struggle with languages having drastically different word orders or grammatical structures, potentially creating suboptimal segmentation points. Furthermore, the alignment quality degrades for distant language pairs or complex sentences with idiomatic expressions and cultural references. While our augmentation strategies help mitigate some issues, they are still constrained by the initial alignment quality.
@@
Appendix A Conversational SimulMT Decoding
@@
5:  $\mathbf{y}_{c}^{\prime},\mathbf{h}^{\prime}\leftarrow\texttt{LLM}(\mathbf{x}_{\text{prompt}},B,\mathbf{h},\text{latest=True})$
@@
Algorithm˜1 presents the details of applying conversational prompts for decoding.